OpenStudy (anonymous):

a sector is inscribed in a square area of 64m^2. find its area. please help :(

OpenStudy (anonymous):

What is a "sector". Be more specific please.

OpenStudy (anonymous):

If we draw a figure of the given problem it will look like as given below  Area of square = side^2 64= side^2 i.e. side of the square = 8 m since side of the square will be the diameter of the sector of the circle therefore the sector of the circle will be semicircle Thus, Area of the sector = area of the semicircle=\[\frac{1}{2}\pi r^2=\frac{1}{2} \times 3.14\times 4^2=\frac{1}{2} \times 3.14\times 16\] \[= 3.14\times 8\] = 25.12 sq m